# Probing For Non-Gravitational Long-Range Dark Matter Interactions

### v2

@@ Abstract
Figure 3: Once per turntable revolution differential acceleration amplitudes along with the corresponding fits to the galactic basis functions. The mean of each data set has been removed. Each two-day long segment was independently fit. Short gaps in the data were due to local construction activity while the long gap between days 275 and 384 was due to a hardware failure. The right panel shows the histogram of the differential acceleration amplitudes, the bottom panel shows the orientations of the pendulum with respect to the vacuum chamber over the data run. The zoom shows a two-day long segment and it’s corresponding fit to the galactic basis functions.
@@
Analysis - The experiment recorded data from July 7, 2024 to July 7, 2025. We rotated the pendulum $180^{\circ}$ with respect to the vacuum chamber on July 31, 2024, February 15, 2025, and May 1, 2025 to reduce systematic effects that may be associated with the pendulum orientation relative to the vacuum chamber (such as tilt and electrostatics). Rotating the pendulum changes the sign of the science signal while maintaining most systematics; these pendulum reorientations were accounted for in the analysis.
The duty cycle of the apparatus was severely restricted due to nearby construction activity and an unexpected hardware failure. Although the data spans 366 days, only 186 days (51%) were free of external disturbances and included in the data set. The measured torsion balance angle was calibrated with a programmed turntable speed change (see Supplementary Material). The angle was converted to differential acceleration by inverting the harmonic oscillator response of the pendulum with a time-domain filter and scaling by the relevant fixed parameters. An additional low-pass filter (1 mHz third-order Butterworth) was applied to minimize the effect of high frequency noise. We corrected for the response of this filter at the turntable frequency (0.4% attenuation, $\sim 54^{\circ}$ phase). The differential acceleration time series was split into two turntable-rotation long sets. Each set was fit to a linear combination of sinusoidal functions with frequencies of harmonics of the turntable angle ($\omega_{TT}$, $2\omega_{TT}$, $3\omega_{TT}$, etc.) and harmonics of the torsional resonance ($\omega_{0}$, $2\omega_{0}$, $3\omega_{0}$, etc.). The full fit function is shown in the Supplementary Material.
@@
The instrument experienced occasional transients due to earthquakes, nearby construction activity and other environmental disturbances. We discarded data sets that fell under two criteria: misfit-squared exceeding the 99th percentile of the corresponding $\chi^{2}$-distribution, and a once per turntable-revolution ($\omega_{TT}$}) acceleration component outside of the 95th percentile of the component’s distribution. We subtracted the mean $\omega_{TT}$}-acceleration amplitude of each data set leaving only the time varying components. Figure 3 shows the surviving once per turntable revolution acceleration amplitudes. The $\omega_{TT}$-acceleration was then extracted in two-day long segments. Segments that had less than a quarter of a sidereal day of data were discarded. Each segment was fit to galactic basis functions. These functions consisted of the location of the center of the Milky Way (RA: $18^{\text{h}}$ Dec: $29^{\circ}$) projected into the horizontal plane of the lab and an orthogonally oriented function. These functions modulate once per sidereal day due to the rotation of the Earth and the location of our lab at $\sim 48^{\circ}$ latitude. The location of the galactic center and the necessary coordinate transformations were computed using the Astropy Astropy Collaboration and Astropy Project Contributors (2022) libraries. This fitting yielded 30 points for the $0^{\circ}$ pendulum orientation and 60 points for $180^{\circ}$ for a total of 90 independent measurements.
@@
$\Delta a_{DM,Be-Al}=0.3\pm 0.6\text{ fm/s}^{2}\quad(1\sigma)$ (1)

### v3

@@ Abstract
Figure 3: Once per turntable revolution differential acceleration amplitudes along with the corresponding fits to the galactic basis functions. The mean of each data set has been removed. Each two-day long segment was independently fit. Short gaps in the data were due to local construction activity. Data for sidereal days 189 to 285 was taken in 2024, while the remaining data was taken in 2025. The right panel shows the histogram of the differential acceleration amplitudes, the bottom panel shows the orientations of the pendulum with respect to the vacuum chamber over the data run. The zoom shows a two-day long segment and it’s corresponding fit to the galactic basis functions.
@@
Analysis - The experiment recorded data from July 7, 2024 to December 25, 2025. We rotated the pendulum $180^{\circ}$ with respect to the vacuum chamber on July 29, 2024, February 15, 2025, May 1, 2025, and September 29, 2025 to reduce systematic effects that may be associated with the pendulum orientation relative to the vacuum chamber (such as tilt and electrostatics). Rotating the pendulum changes the sign of the science signal while maintaining most systematics; these pendulum reorientations were accounted for in the analysis.
The duty cycle of the apparatus was severely restricted due to nearby construction activity and an unexpected hardware failure. Although the data spans 537 days, only 244 days (45%) were free of external disturbances and included in the data set. The measured torsion balance angle was calibrated with a programmed turntable speed change (see Supplementary Material). The angle was converted to differential acceleration by inverting the harmonic oscillator response of the pendulum with a time-domain filter and scaling by the relevant fixed parameters. An additional low-pass filter (1 mHz third-order Butterworth) was applied to minimize the effect of high frequency noise. We corrected for the response of this filter at the turntable frequency (0.4% attenuation, $\sim 54^{\circ}$ phase). The differential acceleration time series was split into two turntable-rotation long sets. Each set was fit to a linear combination of sinusoidal functions with frequencies of harmonics of the turntable angle ($\omega_{TT}$, $2\omega_{TT}$, $3\omega_{TT}$, etc.) and harmonics of the torsional resonance ($\omega_{0}$, $2\omega_{0}$, $3\omega_{0}$, etc.). The full fit function is shown in the Supplementary Material.
@@
The instrument experienced occasional transients due to earthquakes, nearby construction activity and other environmental disturbances. We discarded data sets that fell under two criteria: misfit-squared exceeding the 99th percentile of the corresponding $\chi^{2}$-distribution, and a once per turntable-revolution ($\omega_{TT}$}) acceleration component outside of the 95th percentile of the component’s distribution. We subtracted the mean $\omega_{TT}$}-acceleration amplitude of each data set leaving only the time varying components. Figure 3 shows the surviving once per turntable revolution acceleration amplitudes. The $\omega_{TT}$-acceleration was then extracted in two-day long segments. Segments that had less than a quarter of a sidereal day of data were discarded. Each segment was fit to galactic basis functions. These functions consisted of the location of the center of the Milky Way (RA: $18^{\text{h}}$ Dec: $29^{\circ}$) projected into the horizontal plane of the lab and an orthogonally oriented function. These functions modulate once per sidereal day due to the rotation of the Earth and the location of our lab at $\sim 48^{\circ}$ latitude. The location of the galactic center and the necessary coordinate transformations were computed using the Astropy Astropy Collaboration and Astropy Project Contributors (2022) libraries. This fitting yielded 56 points for the $0^{\circ}$ pendulum orientation and 66 points for $180^{\circ}$ for a total of 122 independent measurements.
@@
$\Delta a_{DM,Be-Al}=0.5\pm 0.6\text{ fm/s}^{2}\quad(1\sigma)$ (1)
A positive $\Delta a$ would correspond to beryllium attracting more towards dark matter. The average acceleration for the 0◦ and 180◦ orientations are individually consistent with zero.